eBook Print item High-Low Method, Cost Formulas The controller of the South Charleston plant of Ravinia, Inc, monitored activities assodated with materials handling costs. The high and low levels of resource usage occurred in September and March for three different resources associated with materials handling. The number of moves is the driver. The total costs of the three resources and the Resource Forklift depreciation: activity output, as measured by moves for the two different levels, are presented as follows: Number of Moves Total Cost Low 5,000 17,000 $3,100 3,100 High Indirect labor: Low High Fuel and oil for forklift: 5,000 $66,000 17,000 128,400 5,000 $3,050 High 17,000 10,370 Required If required, round your answers to two decimal places. Enter a o if required. 1. Determine the cost behavior formula of each resource. Use the high-low method to assess the fixed and variable components. Forklift depreciation: Previous Check My Work 3 more Check My Work uses remaining 816
1. Determine the cost behavior formula of each resource. Use the high-low method to assess the fixed and variable components. Forklift depreclation: Indirect labor: Fuel and oil for forklift: 2. Using your knowledge of cost behavior, predict the cost of each item for an activity output level of 7,000 moves. Forklift depreciation Indirect labor Fuel and oil for forklift 3. Construct a cost formula that can be used to predict the total cost of the three resources combined. If required, round your answers to two decimal Check My Work 3 more Check My Work uses remaining Previous Next 816 PM 2a/01/2019 ^訇届 ENG
eBook Print Itemm Indirect labor: Fuel and oil for forklift: 2. Using your knowledge of cost behavior, predict the cost of each item for an activity output level of 7,000 moves. Forkdift depreciation Indirect labor Fuel and oil for forklift 3. Construct a cost formula that can be used to predict the total cost of the three resources combined. If required, round your answers to two decimal places Materials handling cost . Using this formula, predict the total materials handling cost if activity output is 7,000 moves Previous Next Check My Work 3 more Check My Work uses remaining 816 PM ip
Answer #1

Answer to Requirement 1.

Forklift Depreciation:
As per High-low method:

Variable Cost per Unit = y2 – y1 / x2 – x1

Total Cost at the Highest level of activity (y2) = $3,100
Total Cost at the Lowest level of activity (y1) = $3,100
Units at the highest level of activity (x2) = 17,000
Units at Lowest level of activity (x1) = 5,000

Variable Cost per Unit = ($3,100 - $3,100) / (17,000 – 5,000)
Variable Cost per Unit = $0

Total Cost = Fixed Cost + (Variable Cost * No. of Units) or Total Cost = a + bx
$3,100 = a + ($0 * 5,000)
a = $3,100

Therefore, V = $0, F = $3,100 and Y = $3,100

Indirect Labor:
As per High-low method:

Variable Cost per Unit = y2 – y1 / x2 – x1

Total Cost at the Highest level of activity (y2) = $128,400
Total Cost at the Lowest level of activity (y1) = $66,000
Units at the highest level of activity (x2) = 17,000
Units at Lowest level of activity (x1) = 5,000

Variable Cost per Unit = ($128,400 - $66,000) / (17,000 – 5,000)
Variable Cost per Unit = $5.20

Total Cost = Fixed Cost + (Variable Cost * No. of Units) or Total Cost = a + bx
$66,000 = a + ($5.20 * 5,000)
a = $40,000

Therefore, V = $5.20, F = $40,000 and Y = $40,000 + $5.20X

Fuel and Oil for Forkloft:
As per High-low method:

Variable Cost per Unit = y2 – y1 / x2 – x1

Total Cost at the Highest level of activity (y2) = $10,370
Total Cost at the Lowest level of activity (y1) = $3,050
Units at the highest level of activity (x2) = 17,000
Units at Lowest level of activity (x1) = 5,000

Variable Cost per Unit = ($10,370 - $3,050) / (17,000 – 5,000)
Variable Cost per Unit = $0.61

Total Cost = Fixed Cost + (Variable Cost * No. of Units) or Total Cost = a + bx
$3,050 = a + ($0.61 * 5,000)
a = $0

Therefore, V = $0.61, F = $0 and Y = $0.61X

Answer to Requirement 2.

Forklift Depreciation:
Y = a + bx
Y = $3,100

Therefore, the Cost at an activity level of 7,000 moves is $3,100.

Indirect Labor:
Y = a + bx
Y = $40,000 + $5.20X

Total Cost at an activity level of 7,000 moves = $40,000 + ($5.20 * 7,000)
Total Cost at an activity level of 7,000 moves = $76,400

Fuel and Oil for Forklift:
Y = a + bx
Y = $0.61X

Total Cost at an activity level of 7,000 moves = $0.61 * 7,000
Total Cost at an activity level of 7,000 moves = $4,270

Answer to Requirement 3.

Materials Handling Cost = $3,100 + ($40,000 + $5.20X) + $0.61X
Materials Handling Cost = $43,100 + $5.81X

Answer to Requirement 4.

Total Cost at an activity level of 7,000 moves = $43,100 + ($5.81 * 7,000)
Total Cost at an activity level of 7,000 moves (Y) = $83,770